A cheerleader lifts his 38.6 kg partner straight up off the ground a distance of 0.596 m before releasing her. 2
The acceleration of gravity is 9.8 m/s .
If he does this 28 times, how much work has he done?
1
Expert's answer
2011-10-27T12:48:52-0400
The work of doing that 1 time is:
A1 = F*h = mgh.
Then the total work
is:
A = 28*A1 = 28mgh = 28*38.6*9.8*0.596 = 6312,74 J.